# Scattering of CO from vacant-MoSe2 with O adsorbates: is CO2 formed?

### v1

@@ 1 Introduction
Focusing on the usage of TMDs as catalysts for diverse chemical processes, their applicability is promising. There are various studies exploring the catalytic properties of TMDs for the hydrogen evolution reaction (HER) 31, 32, 33, 24, 3, 34, 25 and the oxygen evolution reaction (OER) 35, 36, 37. These are the two reactions involved in water splitting, a key process for sustainable energy production. In the case of HER, TMD materials appear as good alternatives to precious metals such as platinum. However, pristine TMDs exhibit poor performance for OER due to the large binding energies of intermediate products on the TMD surface 35. Nonetheless, both theory and experiments reveal that the efficiency of these two reactions can be enhanced at the edges of the nanosheets 37, 36, 3, around chalcogen and metal vacancies 31, 32, 24, and by doping the nanosheet 33. Other chemical reactions that are the focus on theoretical studies of the TMD catalytic properties include, for example, CO2 reduction to produce acids and other valuable products 38, 39, 40, 41, 42, 43, 1, 2, 6, N2 reduction to produce ammonia 44, production of ethanol from syngas 45, and also CO oxidation on MoS2 doped with metal atoms 46, 47, on PtTe2 in the presence of O2 48 and on the Janus WSSe surface 49.
@@
In the aforereviewed bibliography, the catalytic properties are commonly determined in terms of the adsorption energies and minimum energy paths of the reaction considered. Being the latter the common procedure, the purpose and novelty of our work are to explore the catalytic properties of MoSe2 for CO2 formation but from a different perspective. In particular, we focus on the dynamics of the recombination process between one O atom adsorbed on the MoSe2 surface with a Se vacancy and a gas-phase CO molecule. This kind of reaction process, which can occur upon a single collision of the gas species with either the adsorbate (direct Eley-Rideal) or the surface (indirect Eley-Rideal) or after few collisions that serve to slow down the incoming atom/molecule (hot atom/molecule abstraction), are expected to efficiently compete with the thermal Langmuir-Hinshelwood recombination when the adsorbate is strongly bound to the surface. This is precisely the case in the dissociative adsorption of O2 on the MoSe2 surface with a Se vacancy. DFT calculations with the revPBE-vdW exchange-correlation functional developed by Dion et al. 50 showed that the dissociative adsorption of O2 is exothermic, being the adsorption energies of the nascent O atoms that adsorb at different sites eV and eV 51. Our new calculations for the recombinative formation of CO2 via an Eley-Rideal mechanism show that the reaction is energetically favored and thus, motivated us to study the process with ab initio molecular dynamics simulations (AIMD). Interestingly, none of the incoming CO molecules is able to abstract an O adatom and form CO2, even if the molecule has enough energy to overcome the entrance barriers for the reaction. These results remark the importance of going beyond static calculations of the reaction energetics in determining the catalyzing character of the material under consideration.
@@
The paper is organized as follows, the general computational settings employed to both model the system and simulate the reaction dynamics with the AIMD methodology are detailed in the Theoretical Methods section. The Results and Discussion section starts with the analysis of the energetics of the main reactions that may occur in the system for the incidence conditions considered, i.e., a gas-phase CO impinging on the MoSe2 surface with a Se vacancy and an oxygen coverage of 0.125 ML. Next, the AIMD results are presented and discussed in the light of the information extracted from the in depth analysis of the trajectories. The summary and conclusions of our work are given in the last section.
@@ 2.1 General DFT computational settings
All DFT calculations are performed with the Vienna ab initio simulation package (vasp) 52 using the revPBE-vdW exchange-correlation functional proposed by Dion et al. 50. The ionic cores are described with the projector augmented wave (PAW) method 53 implemented in VASP 54. Specifically, we use the PAW potentials for Mo, Se, O, and C that have 14, 6, 6, and 4 valence electrons, respectively. The Kohn-Sham states are expanded in a plane-wave basis set using an energy cutoff of 700 eV. The Brillouin zone integration is performed with a -centered 441 Monkhorst–Pack grid of special -points 55. A Gaussian smearing of 0.1 eV is used for electronic state occupancies.
Calculations are performed for the phase of the MoSe2 monolayer with a small concentration of non-interacting Se vacancies (denoted as MoSe2-vSe hereafter). Following ref 51, the latter is adequately described by a 44 surface supercell that contains a single Se vacancy and is separated from its periodic image by 23 Å of vacuum along the surface normal. The value of the hexagonal lattice constant is =3.386 Å 51. The O2 molecule is dissociatively adsorbed in the configuration analyzed in ref 51. Namely, one of the O atoms is adsorbed on top of the Se vacancy and the other one on top of the closest Se atom. The O atom in the vacancy is located at Å below the surface, the O atom on top of the Se atom is located at =1.57 Å above the surface, and the distance between them is d(O-O)=4.01 Å. Dissociation of the O2 molecule in this configuration is exothermic by 3.3 eV. We denote this structure as 2O-MoSe2-vSe.
@@ 2.2 AIMD Initial Conditions
The interaction dynamics of CO with the 2O-MoSe2-vSe surface is studied with total energy-conserving AIMD simulations for an initial substrate temperature of 100 K. The initial positions and velocities of the atoms in the substrate are randomly taken from a canonical ensemble of configurations. The latter is generated by thermalizing the 2O-MoSe2-vSe substrate at 100 K during more than 10 ps using the Nosé-Hoover thermostat 56 implemented in VASP. The use of constant-energy AIMD simulations with an appropriate set of initial positions and velocities constitutes a reasonable approximation to incorporate surface temperature effects during the short simulation time of interest (up to 1.0–1.4 ps in our case) 57. This is a common procedure in gas-surface dynamics simulations 58, 59, 60, 61, 57, 62, 63, 64, 65, 66, 67.
@@
3 Results and Discussion
The calculated reaction energies summarized in table 1 show that both the recombinative CO2 adsorption ( eV) and abstraction ( eV) involving the O atom adsorbed atop Se (O hereafter) are highly exothermic. Recombination with the O atom adsorbed in the Se vacancy (O hereafter) is energetically less favored. In particular, adsorption is exothermic by only eV, while abstraction becomes endothermic, being the reaction energy eV. The different adsorption energy for O ( eV) and O ( eV) 51 is the main reason for the different reaction energies obtained with each adsorbate. CO adsorption near O although exothermic is energetically less favored than CO2 formation. In this adsorption configuration (denoted as CO in table 1), the CO center of mass is practically over a hollow site (see the black-framed panel in Figure 1). Interestingly, the calculated CO adsorption energy on 2O-MoSe2-vSe ( eV) is similar to that of CO adsorbed over a hollow site on pristine MoSe2 ( eV) 51. Nonetheless, since the focus of this work is the recombinative formation of CO2, we did not perform a systematic analysis of all possible CO adsorption configurations on 2O-MoSe2-vSe and, therefore, we cannot discard the existence of CO adsorption configurations more stable than CO.
@@ 3 Results and Discussion
In addition to calculating the reaction energy, we carried out a preliminary inspection of the possible energy barriers the incoming CO may encounter while approaching the surface. For simplicity, we focused on exploring a few paths running close to O. In these calculations the height of the C atom (and the 3 coordinates of one Mo atom located more than 7 Å away the O adsorbates) are always fixed. In some paths, also the coordinates of the C and O atoms forming the molecule are fixed. The rest of degrees of freedom are allowed to relax until forces are smaller than 0.015 eV/Å. The dependence of the potential energy on along these inspection paths is represented in Figure 1. A vertically oriented CO approaching atop O will face an energy barrier of about 0.378 eV, if the CO molecule is forced to stay perpendicular to O by fixing the center of mass coordinates (green diamonds). The abrupt energy decrease at Å corresponds to abstraction of the adsorbed O to form a vertically oriented CO2 (see top and side views of the structures at points and in Figure 1). Note that the energy at is greater than that of desorbed CO2 (asymptotic limit of the CO2 curves plotted by blue and brown triangles in the figure), suggesting that an eventually formed vertical-CO2 will end desorbing. The latter was confirmed with an ulterior structural relaxation of the system configuration at , which ends with the CO2 molecule in vacuum. A similar energy barrier ( eV) is found when only the C coordinates (, , ) are fixed (red squares). In this case CO can tilt while approaching O (see structure in Figure 1) and also ends forming CO2 (see structure ), giving rise to the energy drop observed between points and in the figure. In this case the formed molecule is oriented parallel to the surface and it is energetically more stable than the desorbed CO2. Finally, when only is fixed while approaching over O, CO tilts and is progressively repealed away the initial O–C vertical line, avoiding the energy barriers found in the two previous paths. The minimum energy configuration along this path is shown in the black framed panel. For completeness, we also calculate potential energy curves against for the CO2 formed from CO+O. In these calculations only is fixed. Energies for CO2 approaching over O and atop the Se atom on which O was adsorbed prior recombining are respectively plotted by brown and blue curves in Figure 1. Corresponding minimum energy structures are depicted in the brown and blue framed panels.
@@
Although CO2 formation and desorption upon O abstraction is energetically favored, the previous analysis casts some doubts regarding the likeliness of the reaction. The explored paths present energy barriers in the entrance channel that may difficult the required approach of the CO molecule to O prior recombining. Nevertheless, we cannot exclude the existence of other complex but more favorable reaction paths with lower or nonexistent energy barriers. In this respect, there are many examples showing that the incoming gaseous species can also capture the adsorbate upon reflecting once (indirect ER recombination) or several times (hot atom/molecule recombination) with the surface. Both direct and indirect ER recombination are observed in molecular dynamics (MD) and AIMD simulations of H scattered off Cl/Au(111) 68, 69 and H/Cu(111) 70, 71 and in MD simulations of N scattered off N-covered Ag(111) 72, 73 and N-covered W(110) 74, 75, 76, whereas only indirect ER occurs in the N-covered W(100) 74, 75 even at grazing incidence conditions at which direct ER is likely to occur 76 and, important for the present study, in the formation of CO2 by O atoms impinging on CO/Pt(111) 77. In the case of H impinging on H-covered W(110) and W(100) surfaces, all ER recombinations are indirect 76, being also important the hot atom mechanism as shown in ulterior simulations performed at finite coverage 78, 79, 80.
@@
Considering the rich reactivity observed in other systems, the results of our AIMD calculations are quite surprising. Even if most of the simulated incidence energies ( eV) suffice to overcome the energy barriers explored in Figure 1, none of the trajectories led to CO2 formation. The latter includes both desorption and trapping (adsorption) of the expected CO+O recombination. There are only two possible outcomes observed at the end of simulations that correspond to the impinging CO being either reflected or molecularly trapped on the surface. Processes such as O atomic desorption or O2 recombinative desorption are highly endothermic and, therefore, energetically forbidden for the incidence energies here considered. Figure 2 shows the probabilities for CO reflection and trapping as a function of the incidence energy . At the lowest considered energy ( eV) 65% of the molecules are reflected, whereas a high 35% of them remain dynamically trapped at the end of the simulations (1 ps). Once the first collision with the surface takes place within a short interval of 2–4 fs, the trapped CO molecules experience a second rebound in the vacuum side at very diverse heights ( Å) and are attracted back to the surface. After extending the integration time up to about 1.4 ps, all these molecules stay below 6 Å from the surface, but the final outcome of these trajectories at longer times remains unclear with our limited-time dynamics (see Figure S1 in the Supporting Information). As increases, the trapping probability rapidly decreases in favor of CO reflection. Thus, at the largest eV more than 95% of the molecules are reflected. With very few exceptions, observed at lower incidence energies, the reflection process occurs after the very first collision with the surface. For eV, the first (classical) turning point of the reflected molecules is distributed in three regions depending on whether CO collides near O ( Å ), near O ( Å ), or at other points on the surface ( Å). These turning point regions move closer to the surface and broaden as increases. The turning point distributions for the trapped molecules exhibit a less featured structure. All these observations can be seen in Figures S2, S3, and S4 in the Supporting Information.
@@
A more thorough analysis of the trajectories allows us to better understand the lack of recombination processes. Figures 3 and 4 show the evolution along the incoming trajectory of the CO center of mass coordinates () and the polar angle w.r.t. the surface normal of the CO molecular axis for incidence energies eV and eV, respectively. (Note in passing that the center of mass position at Å provides a clear idea of the area covered by the random distribution over the surface that was used in the initial conditions sampling described in sec AIMD Initial Conditions.) The number of reflected and trapped molecules that reach a certain height during the initial approach is written on each column, providing information on how many molecules have experienced the first collision with the surface between two consecutive height values. The most salient feature in both figures is the inability of the CO molecules to get close enough to O. In the case of eV, already at Å, molecules that were impinging nearby O have been reflected (6 molecules) or have been deviated to remain trapped at the surface (5 molecules). At the higher energy eV (Figure 4), the molecules can get closer. Yet two of the molecules incoming above O have already been reflected at 3.5 Å and at 3 Å all the molecules that were approaching O have been reflected or deviated from the O proximity in the case of the trapped molecules. Interestingly, a majority of the molecules that can get as close to the surface as Å are the ones that are in the region close to O. However, O is buried in the Se topmost layer ( Å) 51 and strongly bound to the substrate, such that recombination does not take place. Regarding the distributions, there is no noticeable effect related to the orientation of the molecules, except for a certain predominance of near parallel orientations () at the closest distances.

### v2

@@ 1 Introduction
Focusing on the usage of TMDs as catalysts for diverse chemical processes, their applicability is promising. There are various studies exploring the catalytic properties of TMDs for the hydrogen evolution reaction (HER) 31, 32, 33, 24, 3, 34, 25 and the oxygen evolution reaction (OER) 35, 36, 37. These are the two reactions involved in water splitting, a key process for sustainable energy production. In the case of HER, TMD materials appear as good alternatives to precious metals such as platinum. However, pristine TMDs exhibit poor performance for OER due to the large binding energies of intermediate products on the TMD surface 35. Nonetheless, both theory and experiments reveal that the efficiency of these two reactions can be enhanced at the edges of the nanosheets 37, 36, 3, around chalcogen and metal vacancies 31, 32, 24, and by doping the nanosheet 33. Other chemical reactions that are the focus on theoretical studies of the TMD catalytic properties include, for example, CO2 reduction to produce acids and other valuable products 38, 39, 40, 41, 42, 43, 1, 2, 6, N2 reduction to produce ammonia 44, production of ethanol from syngas 45, and also CO oxidation on MoS2 doped with metal atoms 46, 47, on PtTe2 in the presence of O2 48, and on the Janus WSSe surface 49.
@@
In the aforereviewed bibliography, the catalytic properties are commonly determined in terms of the adsorption energies and minimum energy paths of the reaction considered. Being the latter the common procedure, the purpose and novelty of our work are to explore the catalytic properties of MoSe2 for CO2 formation but from a different perspective, the reaction dynamics. DFT calculations including van der Waals corrections showed that the dissociative adsorption of O2, which is endothermic by more than 1 eV in the pristine 1H-MoSe2 sheet, becomes highly exothermic in the presence of Se vacancies 50. In light of those results, here we focus on the dynamics of the recombination process between one O atom adsorbed on the MoSe2 surface with a Se vacancy and a gas-phase CO molecule. This kind of reaction process, which can occur upon a single collision of the gas species with either the adsorbate (direct Eley-Rideal) or the surface (indirect Eley-Rideal) or after few collisions that serve to slow down the incoming atom/molecule (hot atom/molecule abstraction), are expected to efficiently compete with the thermal Langmuir-Hinshelwood recombination when the adsorbate is strongly bound to the surface. This is precisely the case in the dissociative adsorption of O2 on the MoSe2 surface with a Se vacancy. As shown in ref 50, the adsorption energies of the nascent O atoms formed upon the dissociative adsorption of O2 in that surface are as strong as eV and eV, depending on the adsorption site. Our new calculations for the recombinative formation of CO2 via an Eley-Rideal mechanism show that the reaction is energetically favored and thus, motivated us to study the process with ab initio molecular dynamics simulations (AIMD). Interestingly, none of the incoming CO molecules is able to abstract an O adatom and form CO2, even if the molecule has enough energy to overcome the entrance barriers for the reaction. These results remark the importance of going beyond static calculations of the reaction energetics in determining the catalyzing character of the material under consideration.
@@
The paper is organized as follows, the general computational settings employed to both model the system and simulate the reaction dynamics with the AIMD methodology are detailed in the Theoretical Methods section. The Results & Discussion section starts with the analysis of the energetics of the main reactions that may occur in the system for the incidence conditions considered, i.e., a gas-phase CO impinging on the MoSe2 surface with a Se vacancy and an oxygen coverage of 0.125 ML. Next, the AIMD results are presented and discussed in light of the information extracted from the in depth analysis of the trajectories. The summary and conclusions of our work are given in the last section.
@@ 2.1 General DFT Computational Settings
All DFT calculations are performed with the Vienna ab initio simulation package (vasp) 51 using the revPBE-vdW exchange-correlation functional proposed by Dion et al. 52. The ionic cores are described with the projector augmented wave (PAW) method 53 implemented in VASP 54. Specifically, we use the PAW potentials for Mo, Se, O, and C that have 14, 6, 6, and 4 valence electrons, respectively. The Kohn-Sham states are expanded in a plane-wave basis set using an energy cutoff of 700 eV. The Brillouin zone integration is performed with a -centered 441 Monkhorst–Pack grid of special -points 55. A Gaussian smearing of 0.1 eV is used for electronic state occupancies.
Calculations are performed for the phase of the MoSe2 monolayer with a small concentration of non-interacting Se vacancies (denoted as MoSe2-vSe hereafter). Following ref 50, the latter is adequately described by a 44 surface supercell that contains a single Se vacancy and is separated from its periodic image by 23 Å of vacuum along the surface normal. The value of the hexagonal lattice constant is =3.386 Å 50. The O2 molecule is dissociatively adsorbed in the configuration analyzed in ref 50. Namely, one of the O atoms is adsorbed on top of the Se vacancy and the other one on top of the closest Se atom. The O atom in the vacancy is located at Å below the surface, the O atom on top of the Se atom is located at =1.57 Å above the surface, and the distance between them is d(O-O)=4.01 Å. Dissociation of the O2 molecule in this configuration is exothermic by 3.3 eV. We denote this structure as 2O-MoSe2-vSe.
Regarding our choice of the revPBE-vdW functional, our previous study of the adsorption properties of different diatomic molecules on MoSe2 and MoSe2-vSe showed that both revPBE-vdW and PBE-D3 56 provide a satisfactory description of the experimental MoSe2 geometrical and electronic structures. Furthermore, even if PBE-D3 tends to overbinding as compared to revPBE-vdW, there were no significant qualitative differences between both functionals in the O, O2, and CO adsorption and dissociation energies and adsorption sites on MoSe2 and MoSe2-vSe.
@@ 2.2 AIMD Initial Conditions
The interaction dynamics of CO with the 2O-MoSe2-vSe surface is studied with total energy-conserving AIMD simulations for an initial substrate temperature of 100 K. The initial positions and velocities of the atoms in the substrate are randomly taken from a canonical ensemble of configurations. The latter is generated by thermalizing the 2O-MoSe2-vSe substrate at 100 K during more than 10 ps using the Nosé-Hoover thermostat 57 implemented in VASP. The use of constant-energy AIMD simulations with an appropriate set of initial positions and velocities constitutes a reasonable approximation to incorporate surface temperature effects during the short simulation time of interest (up to 1.0–1.4 ps in our case) 58. This is a common procedure in gas-surface dynamics simulations 59, 60, 61, 62, 58, 63, 64, 65, 66, 67, 68.
@@
3 Results & Discussion
The calculated reaction energies summarized in table 1 show that both the recombinative CO2 adsorption ( eV) and abstraction ( eV) involving the O atom adsorbed atop Se (O hereafter) are highly exothermic. Recombination with the O atom adsorbed in the Se vacancy (O hereafter) is energetically less favored. In particular, adsorption is exothermic by only eV, while abstraction becomes endothermic, being the reaction energy eV. The different adsorption energy for O ( eV) and O ( eV) 50 is the main reason for the different reaction energies obtained with each adsorbate. CO adsorption near O although exothermic is energetically less favored than CO2 formation. In this adsorption configuration (denoted as CO in table 1), the CO center of mass is practically over a hollow site (see the black-framed panel in Figure 1). Interestingly, the calculated CO adsorption energy on 2O-MoSe2-vSe ( eV) is similar to that of CO adsorbed over a hollow site on pristine MoSe2 ( eV) 50. Nonetheless, since the focus of this work is the recombinative formation of CO2, we did not perform a systematic analysis of all possible CO adsorption configurations on 2O-MoSe2-vSe and, therefore, we cannot discard the existence of CO adsorption configurations more stable than CO.
@@ 3 Results & Discussion
In addition to calculating the reaction energy, we carried out a preliminary inspection of the possible energy barriers the incoming CO may encounter while approaching the surface. For simplicity, we focused on exploring a few paths running close to O. In these calculations the height of the C atom (and the 3 coordinates of one Mo atom located more than 7 Å away the O adsorbates) are always fixed. In some paths, also the coordinates of the C and O atoms forming the molecule are fixed. The rest of degrees of freedom are allowed to relax until forces are smaller than 0.015 eV/Å. The dependence of the potential energy on along these inspection paths is represented in Figure 1. A vertically oriented CO approaching atop O will face an energy barrier of about 0.378 eV, if the CO molecule is forced to stay perpendicular to O by fixing the center of mass coordinates (green diamonds). The abrupt energy decrease at Å corresponds to abstraction of the adsorbed O to form a vertically oriented CO2 (see top and side views of the structures at points a and b in Figure 1). Note that the energy at b is greater than that of desorbed CO2 (asymptotic limit of the CO2 curves plotted by blue and brown triangles in the figure), suggesting that an eventually formed vertical-CO2 will end desorbing. The latter was confirmed with an ulterior structural relaxation of the system configuration at b, which ends with the CO2 molecule in vacuum. A similar energy barrier ( eV) is found when only the C coordinates (, , ) are fixed (red squares). In this case CO can tilt while approaching O (see structure c in Figure 1) and also ends forming CO2 (see structure d), giving rise to the energy drop observed between points c and d in the figure. In this case the formed molecule is oriented parallel to the surface and it is energetically more stable than the desorbed CO2. Finally, when only is fixed while approaching over O, CO tilts and is progressively repealed away the initial O–C vertical line, avoiding the energy barriers found in the two previous paths. The minimum energy configuration along this path is shown in the black framed panel. For completeness, we also calculate potential energy curves against for the CO2 formed from CO+O. In these calculations only is fixed. Energies for CO2 approaching over O and atop the Se atom on which O was adsorbed prior recombining are respectively plotted by brown and blue curves in Figure 1. Corresponding minimum energy structures are depicted in the brown and blue framed panels.
@@
Although CO2 formation and desorption upon O abstraction is energetically favored, the previous analysis casts some doubts regarding the likeliness of the reaction. The explored paths present energy barriers in the entrance channel that may difficult the required approach of the CO molecule to O prior recombining. Nevertheless, we cannot exclude the existence of other complex but more favorable reaction paths with lower or nonexistent energy barriers. In this respect, there are many examples showing that the incoming gaseous species can also capture the adsorbate upon reflecting once (indirect ER recombination) or several times (hot atom/molecule recombination) with the surface. Both direct and indirect ER recombination are observed in molecular dynamics (MD) and AIMD simulations of H scattered off Cl/Au(111) 69, 70 and H/Cu(111) 71, 72 and in MD simulations of N scattered off N-covered Ag(111) 73, 74 and N-covered W(110) 75, 76, 77, whereas only indirect ER occurs in the N-covered W(100) 75, 76 even at grazing incidence conditions at which direct ER is likely to occur 77 and, important for the present study, in the formation of CO2 by O atoms impinging on CO/Pt(111) 78. In the case of H impinging on H-covered W(110) and W(100) surfaces, all ER recombinations are indirect 77, being also important the hot atom mechanism as shown in ulterior simulations performed at finite coverage 79, 80, 81.
@@
Considering the rich reactivity observed in other systems, the results of our AIMD calculations are quite surprising. Even if most of the simulated incidence energies ( eV) suffice to overcome the energy barriers explored in Figure 1, none of the trajectories led to CO2 formation. The latter includes both desorption and trapping (adsorption) of the expected CO+O recombination. There are only two possible outcomes observed at the end of simulations that correspond to the impinging CO being either reflected or molecularly trapped on the surface. Processes such as O atomic desorption or O2 recombinative desorption are highly endothermic and, therefore, energetically forbidden for the incidence energies here considered. Figure 2 shows the probabilities for CO reflection and trapping as a function of the incidence energy . At the lowest considered energy ( eV) 65% of the molecules are reflected, whereas a high 35% of them remain dynamically trapped at the end of the simulations (1 ps). Once the first collision with the surface takes place within a short interval of 200–400 fs, the trapped CO molecules experience a second rebound in the vacuum side at very diverse heights ( Å) and are attracted back to the surface. After extending the integration time up to about 1.4 ps, all these molecules stay below 6 Å from the surface, but the final outcome of these trajectories at longer times remains unclear with our limited-time dynamics (see Figure S1 in the Supporting Information). As increases, the trapping probability rapidly decreases in favor of CO reflection. Thus, at the largest eV more than 95% of the molecules are reflected. With very few exceptions, observed at lower incidence energies, the reflection process occurs after the very first collision with the surface. For eV, the first (classical) turning point of the reflected molecules is distributed in three regions depending on whether CO collides near O ( Å ), near O ( Å ), or at other points on the surface ( Å). These turning point regions move closer to the surface and broaden as increases. The turning point distributions for the trapped molecules exhibit a less featured structure. All these observations can be seen in Figures S2, S3, and S4 in the Supporting Information.
@@
A more thorough analysis of the trajectories allows us to better understand the lack of recombination processes. Figures 3 and 4 show the evolution along the incoming trajectory of the CO center of mass coordinates () and the polar angle w.r.t. the surface normal of the CO molecular axis for incidence energies eV and eV, respectively. (Note in passing that the center of mass position at Å provides a clear idea of the area covered by the random distribution over the surface that was used in the initial conditions sampling described in sec AIMD Initial Conditions.) The number of reflected and trapped molecules that reach a certain height during the initial approach is written on each column, providing information on how many molecules have experienced the first collision with the surface between two consecutive height values. The most salient feature in both figures is the inability of the CO molecules to get close enough to O. In the case of eV, already at Å, molecules that were impinging nearby O have been reflected (6 molecules) or have been deviated to remain trapped at the surface (5 molecules). At the higher energy eV (Figure 4), the molecules can get closer. Yet two of the molecules incoming above O have already been reflected at 3.5 Å and at 3 Å all the molecules that were approaching O have been reflected or deviated from the O proximity in the case of the trapped molecules. Interestingly, a majority of the molecules that can get as close to the surface as Å are the ones that are in the region close to O. However, O is buried in the Se topmost layer ( Å) 50 and strongly bound to the substrate, such that recombination does not take place. Regarding the distributions, there is no noticeable effect related to the orientation of the molecules, except for a certain predominance of near parallel orientations () at the closest distances.